OpenStudy (anonymous):

Find all solutions x in the interval [0,2pi) cos^2 (x/2) + 2cos (x/2) = -1

OpenStudy (tkhunny):

Have you considered simply factoring? Can you solve this? \(x^{2} + 2x + 1 = 0\)?

OpenStudy (anonymous):

Yes (x+1)^2=0

OpenStudy (tkhunny):

Excellent. Do EXACTLY the same thing with your original equation.

OpenStudy (anonymous):

Hmm.. Would it be cos(x/2)(cos(x/2)+2)=-1 ? Or am I wrong..

OpenStudy (tkhunny):

That was not EXACTLY the same thing. Add 1 to both sides and give it another go. Remember, there is virtually no point to factoring unless one side of the equation is ZERO (0).

OpenStudy (anonymous):

Okay.. so by adding 1 to each side, my new problem would be cos^2(x/2)+2cos(x/2)+1=0 .. and once factoring that, it would become 1/2 (4cos(x/2)+cos(x)+3) = 0 I still feel like that may be wrong..

OpenStudy (anonymous):

I'm sorry. I'm not understanding this at all..

OpenStudy (tkhunny):

A few quick examples, starting with the one YOU already solved. (1) \(x^{2} + 2x + 1 = 0\) \((x+1)^{2} = 0\) (2) \(z^{2} + 2z + 1 = 0\) \((z+1)^{2} = 0\) (3) \(Frog^{2} + 2Frog + 1 = 0\) \((Frog+1)^{2} = 0\) (4) \(\sin^{2}(x) + 2\sin(x) + 1 = 0\) \((\sin(x)+1)^{2} = 0\) (5) \((Eiffel\;Tower)^{2} + 2(Eiffel\;Tower) + 1 = 0\) \(((Eiffel\;Tower)+1)^{2} = 0\) (6) \(\cos^{2}(x/2) + 2\cos(x/2) + 1 = 0\) \((\cos(x/2)+1)^{2} = 0\) EXACTLY like you did my little exercise. EXACTLY.

OpenStudy (anonymous):

Oh! Okay, I see that. Didn't realize it would be that simple.. Now that is factored though, how do I find the solutions of x in the interval [0,2pi) ?

OpenStudy (tkhunny):

Again, EXACTLY like you would with the one that you solved. \((x+1)^{2} = 0\) then \(x+1 = 0\;and\;x = -1\) Let's see what you get.